A 1.0-mm diameter extension cord is made of metal having a resistivity of 1.68 × 10-8 Ω • m. When it carries a current of 15 A, what is the potential difference between two points in the cord that are 100 m apart?
Answer #1

dia=1mm=1*10^-3m

p=1.68 × 10-8 Ω • m.

I=15 amp

d=100m

A=pie*(10^-30^2/4=7.85*10^-7 m^2

R=PL/A

=1.68 × 10-8 *100/7.85*10^-7

=2.14 ohm

V=IR

=15*2.14

=32V